# Function spaces for decoupling

### v2

@@ 1.1. Setting
for $1\leq p\leq\infty$. These embeddings allow one to recover the optimal $L^{p}(\mathbb{R}^{n})$ regularity of Fourier integral operators due to Seeger, Sogge and Stein [34], and the accompanying sharp fixed-time $L^{p}(\mathbb{R}^{n})$ regularity for wave equations. Moreover, the Hardy spaces for Fourier integral operators have been applied to wave equations with rough coefficients [15, 20] and to nonlinear wave equations [30, 27] in ways that would not be possible by working on $L^{p}(\mathbb{R}^{n})$ directly.
@@
However, this development only concerns applications of decoupling inequalities to wave equations, and it shows that the Hardy spaces for Fourier integral operators are natural spaces of initial data in that regard. In the present article, we develop the connection between decoupling inequalities and invariant spaces for wave propagators further, by introducing function spaces that capture the decoupling inequalities for the sphere and the light cone in full generality. We subsequently use these function spaces to obtain improved fractional integration theorems, local smoothing estimates, and well-posedness results for nonlinear wave equations.
@@ 1.2. Main results
Let $p,q\in[1,\infty)$ and $s\in\mathbb{R}$. Then $e^{it\sqrt{-\Delta}}:\mathcal{L}_{W,s}^{q,p}(\mathbb{R}^{n})\to\mathcal{L}_{W,s}^{q,p}(\mathbb{R}^{n})$ is bounded for all $t\in\mathbb{R}$.
On the other hand, there exists a compactly supported Fourier integral operator $T$ such that $T:\mathcal{L}_{W,0}^{q,p}(\mathbb{R}^{n})\to\mathcal{L}_{W,0}^{q,p}(\mathbb{R}^{n})$ is not bounded if $p\neq q$.
@@
More precisely, for each $R\geq 2$, let $V_{R}\subseteq S^{n-1}$ be a maximal collection of unit vectors satisfying $|\nu-\nu^{\prime}|\geq R^{-1/2}$ for all $\nu,\nu^{\prime}\in V_{R}$, and let $(\chi_{\nu})_{\nu\in V_{R}}$ be an associated partition of unity of functions homogeneous of degree 00 (see Section 7.1 for details). Set
@@ 1.5. Notation and terminology
The natural numbers are =${\mathbb{N}}=\{1,2,\ldots\}$, and +:=$\mathbb{Z}_{+}:={\mathbb{N}}\cup\{0\}$. We write $\mathbb{R}_{+}$ for $(0,\infty)$, endowed with the Haar measure $\mathrm{d}\sigma/\sigma$. Moreover, $S^{n-1}$ is the unit sphere in $\mathbb{R}^{n}$, with the unit normalized surface measure $\mathrm{d}\omega$. Throughout this article we fix $n\in{\mathbb{N}}$ with $n\geq 2$.
@@ 2.1. A family of norms
where $\Pi_{\omega}^{\perp}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ is the orthogonal projection onto the complement of the span of $\omega$. Note that, with $P_{\omega}(x):=x+(\omega\cdot x)\omega$, one has
@@
Let $|x|_{\omega}$ be the unique $\sigma_{x}>0$ such that $|A_{\omega,1/\sigma_{x}}(x)|=1$. Then $|\cdot|_{\omega}$ is a norm on $\mathbb{R}^{n}$, in the sense of [37, XIII.5] (see [6, Section 1.4]), and $|A_{\omega,\sigma}x|_{\omega}=\sigma|x|_{\omega}$ for all $\sigma>0$. Let $\mathbb{R}^{n}_{\omega}$ be the metric measure space obtained by endowing $\mathbb{R}^{n}$ with the associated metric and the standard Lebesgue measure. For $\tau>0$, we write
@@
It follows from Lemma 2.1 that the metric associated with $|\cdot|_{\omega}$ is essentially the restriction to the subspace $\mathbb{R}^{n}\times\{\omega\}\subseteq S^{*}\mathbb{R}^{n}$ of the metric on the cosphere bundle from [19, 35]. Similarly, the square of the metric associated with $|\cdot|_{\omega}$ essentially coincides with the quasi-distance from [15].
@@
For a fixed $\omega\in S^{n-1}$, the statement is a special case of [37, Sections II.1 and II.5.14] or [32, Lemma 2.8]. But by rotation, the resulting constant is independent of $\omega$.∎
@@
with the first inequality being classical (see e.g. [37, Sections II.2.1 and II.5.14]). Then Lemma 2.3 concludes the proof. ∎
In the next section, we will combine Lemma 2.3 with the following pointwise variant of [32, Lemma 2.3].
@@
One uses the mean value theorem and the fact that balls in $\mathbb{R}^{n}_{\omega}$ are convex with respect to the Euclidean metric. For a proof of (2.9) in the case where $\rho=1$, see e.g. [39, Theorem 1.3.1]. ∎
@@ 2.2. Hardy spaces
Then $\Psi_{\omega}\in C^{\infty}_{c}(\mathbb{R}^{n})$, due to the fact that $\xi\mapsto|\xi|_{\omega}$ is smooth on $\mathbb{R}^{n}\setminus\{0\}$ (see [6, Lemma 1.5]). Moreover, $\mathrm{supp}(\Psi_{\omega})\subseteq\{\xi\in\mathbb{R}^{n}\mid|\xi|_{\omega}\in[1/2,2]\}$, and
@@
We will write $\Psi_{\omega}(A_{\omega,\sigma}D)$ for the Fourier multiplier with symbol $\xi\mapsto\Psi_{\omega}(A_{\omega,\sigma}\xi)$. Note that $\Psi_{\omega}(A_{\omega,\sigma}D)$ has kernel $\sigma^{-(n+1)}\mathcal{F}^{-1}(\Psi_{\omega})(A_{\omega,1/\sigma}y)$ for $y\in\mathbb{R}^{n}$. Moreover, for each $N\geq 0$ there exists a $C_{N}\geq 0$, independent of $\omega$, $\sigma$ and $y$, such that
@@
As was the case for $H^{p}(\mathbb{R}^{n})$ and $\mathcal{H}^{p}(\mathbb{R}^{n})$, for all $w\in S^{n-1}$ one has
@@
with equivalence of norms (see [7, Theorem 1.2]). Moreover, $H^{1}_{\omega}(\mathbb{R}^{n})\subseteq L^{1}(\mathbb{R}^{n})$ continuously, by [14, Theorem 2.7].
@@
On the other hand, the analogue of Lemma 2.3 also holds for the Littlewood–Paley maximal function $M$ (see [37, Section II.1]). Combined with (2.22), this yields
@@
The first result concerns the anisotropic fractional integration operator $I_{\omega,s}$, given for $\omega\in S^{n-1}$, $s>0$, $f\in\mathcal{S}(\mathbb{R}^{n})$ with $0\notin\mathrm{supp}(\widehat{f}\,)$, and $x\in\mathbb{R}^{n}$, by
@@
This allows us to use the results on singular integrals on homogeneous groups, as detailed in, for example, [14] or [37, Chapter XIII, Section 5]. More precisely, we consider $\mathbb{R}^{n}$ as a Lie group with the usual addition structure, and let
@@
with $a_{1}=\dots=a_{n-1}=1$ and $a_{n}=2$ (see Example 5.2.1 in [37, Chapter XIII, Section 5]). Define a corresponding norm function $\rho(x):=\max_{1\leq j\leq n}|x_{j}|^{1/a_{j}}$. If $K$ is the tempered distribution given by the inverse Fourier transform of $m$, then $K$ satisfies equation (61) in [37, Chapter XIII, Section 5], namely
@@
where $a=a_{1}+\dots+a_{n}=n+1$ (one can verify this by a simple adaptation of the proof of Proposition 2 in [37, Chapter IV, Section 4.4]). In the terminology of [14, p.185], this says that $K$ is a kernel of type $(0,r)$ for any positive integer $r$. Thus the case $\alpha=0$ of [14, Theorem 6.10] guarantees that $m(D)$ is bounded on $H^{p}_{\omega}(\mathbb{R}^{n})$ for all $0<p<\infty$; we only need the case $1\leq p<\infty$ here. (For $1<p<\infty$, one can also use Theorem 3 of [37, Chapter I, Section 5] together with duality, as pointed out at the beginning of [37, Chapter XIII, Section 5.3.1]). ∎
@@ 2.3. Fourier integral operators
    $\sup_{(x,\eta)\in\mathbb{R}^{2n}\setminus o}|\partial_{x}^{\beta}\partial_{\eta}^{\alpha}\Phi(x,\hat{\eta})|<\infty$ for all $\alpha,\beta\in\mathbb{Z}_{+}^{n}$ with $|\alpha|+|\beta|\geq 2$;
@@
    $(\partial_{\eta}\Phi(x,\eta),\eta)\mapsto(x,\partial_{x}\Phi(x,\eta))$ is a well-defined bijection on $\mathbb{R}^{2n}\setminus o$,
@@
If (3) holds, then $(\partial_{\eta}\Phi(x,\eta),\eta)\mapsto(x,\partial_{x}\Phi(x,\eta))$ is a homogeneous canonical transformation on $\mathbb{R}^{2n}\setminus o$, and the canonical relation of $T$ is the graph of this transformation.
If (1) and (2) hold, then (3) holds if and only if $\eta\mapsto\partial_{x}\Phi(x,\eta)$ is a bijection on $\mathbb{R}^{n}\setminus\{0\}$ for each $x\in\mathbb{R}^{n}$, by Hadamard’s global inverse function theorem [26, Theorem 6.2.8]. Another application of the global inverse function theorem then shows that condition (3) is superfluous for $n\geq 3$. Moreover, (3) holds if $\Phi(x,\eta)=x\cdot\eta+\phi(\eta)$ for some $\phi\in C^{\infty}(\mathbb{R}^{n}\setminus\{0\})$ which is positively homogeneous of degree $1$. This case will be considered frequently by us, and it is characterized by the property that $\partial_{x}\Phi(x,\eta)=\eta$ for all $(x,\eta)\in\mathbb{R}^{2n}\setminus o$.
Recall that a compactly supported Fourier integral operator of order $m$ and type $(1,0)$, associated with a local canonical graph, can, modulo an operator with a Schwartz kernel which is a Schwartz function, be expressed as a finite sum of operators which in appropriate coordinate systems are Fourier integral operators in standard form (see e.g. [36, Proposition 6.2.4]), with symbols in $S^{m}_{1,0}$. In this case, the symbol $a$ has compact support in the $x$-variable, and thus (1) and (2) hold on the support of $a$, while the map in (3) is a locally well-defined homogeneous canonical transformation. By contrast, for operators associated with a global canonical graph as in Definition 2.12, the symbols are not required to have compact spatial support, but the conditions on the phase function hold uniformly on all of $\mathbb{R}^{2n}\setminus o$.
@@ 3.1. Wave packets
    The map $(\xi,\omega)\mapsto\varphi_{\omega}(\xi)$ is measurable on $\mathbb{R}^{n}\times S^{n-1}$, and
@@
For the construction of such a collection, see e.g. [28, Section 3.1]. In fact, the functions constructed there have slightly different support properties than in (1). However, the construction in [28, Section 3.1] also yields functions satisfying (1), if one shrinks the support of the function $\varphi$ used there, and slightly modifies the definition of $\varphi_{\omega}$ itself. Moreover, instead of (3.1), from [28, Remark 3.3] one obtains
@@
for a standard symbol $m\in S^{0}(\mathbb{R}^{n})$. However, the proof of [28, Lemma 3.2] also shows that $\sqrt{m}\in S^{0}(\mathbb{R}^{n})$, so that one may replace $\varphi_{\omega}$ by $\sqrt{m}\varphi_{\omega}$ to arrive at (3.1).
@@
(3.2) $\mathrm{supp}(\varphi_{\omega})\subseteq\{\xi\in\mathbb{R}^{n}\mid\tfrac{1}{4}\leq|\xi|_{\omega},\tfrac{1}{\sqrt{5}}|\xi|^{1/2}\leq|\xi|_{\omega}\leq 2|\xi|^{1/2}\}$
@@
We may assume that $\rho_{0}\in C^{\infty}_{c}(\mathbb{R}^{n})$, as in [19, Section 4].
These functions have the following properties, similar to those in [19, Lemma 4.1].
@@
$\tfrac{1}{\sqrt{10}}\sigma^{-1/2}\leq|\xi|_{\omega}\leq\sqrt{8}\sigma^{-1/2}.$
@@
The proof is completely analogous to that of [19, Lemma 4.1]. The first few and the last statement follow from properties (1), (2) and (3) of $(\varphi_{\omega})_{\omega\in S^{n-1}}$, together with (3.2) and the properties of $\Psi$. One then obtains (3.5) from integration by parts, using also the equivalent expression for $|\cdot|_{\omega}$ from Lemma 2.1.∎

### v3

@@ 1.1. Setting
for $1\leq p\leq\infty$. These embeddings allow one to recover the optimal $L^{p}(\mathbb{R}^{n})$ regularity of Fourier integral operators due to Seeger, Sogge and Stein [36], and the accompanying sharp fixed-time $L^{p}(\mathbb{R}^{n})$ regularity for wave equations. Moreover, the Hardy spaces for Fourier integral operators have been applied to wave equations with rough coefficients [16, 21] and to nonlinear wave equations [32, 29] in ways that do not seem possible when working on $L^{p}(\mathbb{R}^{n})$ directly.
@@
However, this development only concerns applications of decoupling inequalities to wave equations, and it shows that the Hardy spaces for Fourier integral operators are natural spaces of initial data in that regard. In the present article, we develop the connection between decoupling inequalities and invariant spaces for wave propagators further, by introducing function spaces that capture the decoupling inequalities for the sphere and the light cone in full generality. In particular, these are not just suitable spaces of initial data for the Euclidean wave equation, but also for its solutions. Moreover, by working with general $\ell^{q}$ decoupling inequalities instead of merely considering the case where $q=p$, we obtain improved fractional integration theorems, as well as new local smoothing estimates and well-posedness results for nonlinear wave equations.
@@ 1.2. Main results
Let $p,q\in[1,\infty)$ and $s\in\mathbb{R}$. Then $e^{it\sqrt{-\Delta}}:\mathcal{L}_{W,s}^{q,p}(\mathbb{R}^{n})\rightarrow\mathcal{L}_{W,s}^{q,p}(\mathbb{R}^{n})$ is bounded for all $t\in\mathbb{R}$.
On the other hand, there exists a compactly supported Fourier integral operator $T$ such that $T:\mathcal{L}_{W,0}^{q,p}(\mathbb{R}^{n})\rightarrow\mathcal{L}_{W,0}^{q,p}(\mathbb{R}^{n})$ is not bounded if $p\neq q$.
@@
More precisely, for each $R\geq 2$, let $V_{R}\subseteq S^{n-1}$ be a maximal collection of unit vectors satisfying $|\nu-\nu^{\prime}|\geq R^{-1/2}$ for all $\nu,\nu^{\prime}\in V_{R}$, and let $(\chi_{\nu})_{\nu\in V_{R}}$ be an associated partition of unity of functions homogeneous of degree 0 (see Section 7.1 for details). Set
@@ 1.5. Notation and terminology
The natural numbers are ${\mathbb{N}}=\{1,2,\ldots\}$, and $\mathbb{Z}_{+}:={\mathbb{N}}\cup\{0\}$. We write $\mathbb{R}_{+}$ for $(0,\infty)$, endowed with the Haar measure $\mathrm{d}\sigma/\sigma$. Moreover, $S^{n-1}$ is the unit sphere in $\mathbb{R}^{n}$, with the unit normalized surface measure $\mathrm{d}\omega$. Throughout this article we fix $n\in{\mathbb{N}}$ with $n\geq 2$.
@@ 2.1. A family of norms
where $\Pi_{\omega}^{\perp}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ is the orthogonal projection onto the complement of the span of $\omega$. Note that, with $P_{\omega}(x):=x+(\omega\cdot x)\omega$, one has
@@
Let $|x|_{\omega}$ be the unique $\sigma_{x}>0$ such that $|A_{\omega,1/\sigma_{x}}(x)|=1$. Then $|\cdot|_{\omega}$ is a norm on $\mathbb{R}^{n}$, in the sense of [39, XIII.5] (see [6, Section 1.4]), and $|A_{\omega,\sigma}x|_{\omega}=\sigma|x|_{\omega}$ for all $\sigma>0$. Let $\mathbb{R}^{n}_{\omega}$ be the metric measure space obtained by endowing $\mathbb{R}^{n}$ with the associated metric and the standard Lebesgue measure. For $\tau>0$, we write
@@
It follows from Lemma 2.1 that the metric associated with $|\cdot|_{\omega}$ is, at small distances, essentially the restriction to the subspace $\mathbb{R}^{n}\times\{\omega\}\subseteq S^{*}\mathbb{R}^{n}$ of the metric on the cosphere bundle from [20, 37]. Similarly, the square of the metric associated with $|\cdot|_{\omega}$ essentially coincides with the quasi-distance from [16].
@@
For a fixed $\omega\in S^{n-1}$, the statement is a special case of [39, Sections II.1 and II.5.14] or [34, Lemma 2.8]. But by rotation, the resulting constant is independent of $\omega$.∎
@@
with the first inequality being classical (see e.g. [39, Sections II.2.1 and II.5.14]). Then Lemma 2.3 concludes the proof. ∎
In the next section, we will combine Lemma 2.3 with the following pointwise variant of [34, Lemma 2.3].
@@
One uses the mean value theorem and the fact that balls in $\mathbb{R}^{n}_{\omega}$ are convex with respect to the Euclidean metric. For a proof of (2.9) in the case where $\rho=1$, see e.g. [42, Theorem 1.3.1]. ∎
@@ 2.2. Hardy spaces
Then $\Psi_{\omega}\in C^{\infty}_{c}(\mathbb{R}^{n})$, due to the fact that $\xi\mapstochar\rightarrow|\xi|_{\omega}$ is smooth on $\mathbb{R}^{n}\setminus\{0\}$ (see [6, Lemma 1.5]). Moreover, $\mathrm{supp}(\Psi_{\omega})\subseteq\{\xi\in\mathbb{R}^{n}\mid|\xi|_{\omega}\in[1/2,2]\}$, and
@@
We will write $\Psi_{\omega}(A_{\omega,\sigma}D)$ for the Fourier multiplier with symbol $\xi\mapstochar\rightarrow\Psi_{\omega}(A_{\omega,\sigma}\xi)$. Note that $\Psi_{\omega}(A_{\omega,\sigma}D)$ has kernel $\sigma^{-(n+1)}\mathcal{F}^{-1}(\Psi_{\omega})(A_{\omega,1/\sigma}y)$ for $y\in\mathbb{R}^{n}$. Moreover, for each $N\geq 0$ there exists a $C_{N}\geq 0$, independent of $\omega$, $\sigma$ and $y$, such that
@@
As was the case for $H^{p}(\mathbb{R}^{n})$ and $\mathcal{H}^{p}(\mathbb{R}^{n})$, for all $\omega\in S^{n-1}$ one has
@@
with equivalence of norms (see [7, Theorem 1.2]). Moreover, $H^{1}_{\omega}(\mathbb{R}^{n})\subseteq L^{1}(\mathbb{R}^{n})$ continuously, by [15, Theorem 2.7].
@@
On the other hand, the analogue of Lemma 2.3 also holds for the Hardy–Littlewood maximal function $M$ (see [39, Section II.1]). Combined with (2.22), this yields
@@
[…]
@@
This allows us to use the results on singular integrals on homogeneous groups, as detailed in, for example, [15] or [39, Chapter XIII, Section 5]. More precisely, we consider $\mathbb{R}^{n}$ as a Lie group with the usual addition structure, and let
@@
with $a_{1}=\dots=a_{n-1}=1$ and $a_{n}=2$ (see Example 5.2.1 in [39, Chapter XIII, Section 5]). Define a corresponding norm function $\rho(x):=\max_{1\leq j\leq n}|x_{j}|^{1/a_{j}}$. If $K$ is the tempered distribution given by the inverse Fourier transform of $m$, then $K$ satisfies equation (61) in [39, Chapter XIII, Section 5], namely
@@
where $a=a_{1}+\dots+a_{n}=n+1$ (one can verify this by a simple adaptation of the proof of Proposition 2 in [39, Chapter IV, Section 4.4]). In the terminology of [15, p.185], this says that $K$ is a kernel of type $(0,r)$ for any positive integer $r$. Thus the case $\alpha=0$ of [15, Theorem 6.10] guarantees that $m(D)$ is bounded on $H^{p}_{\omega}(\mathbb{R}^{n})$ for all $0<p<\infty$; we only need the case $1\leq p<\infty$ here. (For $1<p<\infty$, one can also use Theorem 3 of [39, Chapter I, Section 5] together with duality, as pointed out at the beginning of [39, Chapter XIII, Section 5.3.1]). ∎
@@ 2.3. Fourier integral operators
[…]
@@
    $(\partial_{\eta}\Phi(x,\eta),\eta)\mapstochar\rightarrow(x,\partial_{x}\Phi(x,\eta))$ is a well-defined bijection on $\mathbb{R}^{2n}\setminus o$,
@@
If (3) holds, then $(\partial_{\eta}\Phi(x,\eta),\eta)\mapstochar\rightarrow(x,\partial_{x}\Phi(x,\eta))$ is a homogeneous canonical transformation on $\mathbb{R}^{2n}\setminus o$, and the canonical relation of $T$ is the graph of this transformation.
If (1) and (2) hold, then (3) holds if and only if $\eta\mapstochar\rightarrow\partial_{x}\Phi(x,\eta)$ is a bijection on $\mathbb{R}^{n}\setminus\{0\}$ for each $x\in\mathbb{R}^{n}$, by Hadamard’s global inverse function theorem [28, Theorem 6.2.8]. Another application of the global inverse function theorem then shows that condition (3) is superfluous for $n\geq 3$. Moreover, (3) holds if $\Phi(x,\eta)=x\cdot\eta+\phi(\eta)$ for some $\phi\in C^{\infty}(\mathbb{R}^{n}\setminus\{0\})$ which is positively homogeneous of degree $1$. This case will be considered frequently by us, and it is characterized by the property that $\partial_{x}\Phi(x,\eta)=\eta$ for all $(x,\eta)\in\mathbb{R}^{2n}\setminus o$.
Recall that a compactly supported Fourier integral operator of order $m$ and type $(1,0)$, associated with a local canonical graph, can, modulo an operator with a Schwartz kernel which is a Schwartz function, be expressed as a finite sum of operators which in appropriate coordinate systems are Fourier integral operators in standard form (see e.g. [38, Proposition 6.2.4]), with symbols in $S^{m}_{1,0}$. In this case, the symbol $a$ has compact support in the $x$-variable, and thus (1) and (2) hold on the support of $a$, while the map in (3) is a locally well-defined homogeneous canonical transformation. By contrast, for operators associated with a global canonical graph as in Definition 2.12, the symbols are not required to have compact spatial support, but the conditions on the phase function hold uniformly on all of $\mathbb{R}^{2n}\setminus o$.
@@ 3.1. Wave packets
    The map $(\xi,\omega)\mapstochar\rightarrow\varphi_{\omega}(\xi)$ is measurable on $\mathbb{R}^{n}\times S^{n-1}$, and
@@
For the construction of such a collection, see e.g. [30, Section 3.1]. In fact, the functions constructed there have slightly different support properties than in (1). However, the construction in [30, Section 3.1] also yields functions satisfying (1), if one shrinks the support of the function $\varphi$ used there, and slightly modifies the definition of $\varphi_{\omega}$ itself. Moreover, instead of (3.1), from [30, Remark 3.3] one obtains
@@
for a standard symbol $m\in S^{0}(\mathbb{R}^{n})$. However, the proof of [30, Lemma 3.2] also shows that $\sqrt{m}\in S^{0}(\mathbb{R}^{n})$, so that one may replace $\varphi_{\omega}$ by $\sqrt{m}\varphi_{\omega}$ to arrive at (3.1).
@@
(3.2) $\mathrm{supp}(\varphi_{\omega})\subseteq\{\xi\in\mathbb{R}^{n}\mid\tfrac{1}{4}\leq|\xi|_{\omega},\tfrac{1}{2\sqrt{5}}|\xi|^{1/2}\leq|\xi|_{\omega}\leq 2|\xi|^{1/2}\}$
@@
We may assume that $\rho_{0}\in C^{\infty}_{c}(\mathbb{R}^{n})$, as in [20, Section 4].
These functions have the following properties, similar to those in [20, Lemma 4.1].
@@
$\tfrac{1}{2\sqrt{10}}\sigma^{-1/2}\leq|\xi|_{\omega}\leq\sqrt{8}\sigma^{-1/2}.$
@@
The proof is completely analogous to that of [20, Lemma 4.1]. The first few and the last statement follow from properties (1), (2) and (3) of $(\varphi_{\omega})_{\omega\in S^{n-1}}$, together with (3.2) and the properties of $\Psi$. One then obtains (3.5) from integration by parts, using also the equivalent expression for $|\cdot|_{\omega}$ from Lemma 2.1.∎